# Over-The-Air Phase Calibration of Spaceborne Phased Array for LEO Satellite Communications

### v1

@@ III Over-The-Air Phases Calibration Algorithm
In this section, we propose a block coordinate descent algorithm to solve (6): we first initialize $\bf\Omega$ with all entries being 11, and estimate the channel $\bf H$}, then given the estimated $\bf H$} and update $\bf\Omega$; as the above iteration goes on, the final result can be obtained until (6) converges. Given the estimated $\bf\Omega$ and $\bf H$}, we also point out the phase ambiguity between them, and remove it by pre-assigning values to the corresponding variables.
@@ IV-A Cramér Rao Bounds
Considering the phase ambiguity, we collect the rest phases of elements of $\bf\Omega$ into the vector $\bf p$}, where ${\bf p}=\left[\angle{\hbox{$\bf\Omega$}_{2,1}},\dots,\angle{\hbox{$\bf\Omega$}_{M_{t},1}},\dots,\angle{\hbox{$\bf\Omega$}_{1,2}},\dots,\angle{\hbox{$\bf\Omega$}_{M_{t}-1,2}},\angle{\hbox{$\bf\Omega$}_{M_{t},N_{\rm RF}}},\right$. $\left.\dots,\angle{\hbox{$\bf\Omega$}_{1,N_{\rm RF}}},\dots,\angle{\hbox{$\bf\Omega$}_{M_{t}-1,N_{\rm RF}}},\angle{\hbox{$\bf\Omega$}_{M_{t},N_{\rm RF}}}\right]^{T}$. Hence for the calibration problem, the unknown parameters can be constructed into $\mbox{$\eta$}=\left[{\bf p}^{T},\theta_{r},\phi_{r},{\rm Real}(\gamma),{\rm Imag}(\gamma)\right]^{T}$. To derive the CRB for the phase calibration problem, we have from (17) that $\bar{{\bf y}}={\bf D}{\mbox{$\omega$}}+\bar{{\bf z}}$, where ${\bf D}=\mathop{\rm blkdiag}\{{\bf C}_{1},{\bf C}_{2},\dots,{\bf C}_{N_{\rm RF}}\}$, $\bar{{\bf y}}=[\bar{{\bf y}}_{1}^{T},\bar{{\bf y}}_{2}^{T},\dots,\bar{{\bf y}}_{N_{\rm RF}}^{T}]^{T}$, ${\mbox{$\omega$}}=[\mbox{$\omega$}_{1}^{T},\mbox{$\omega$}_{2}^{T},\dots,\mbox{$\omega$}_{N_{\rm RF}}^{T}]^{T}$, and $\bar{{\bf z}}=[\bar{{\bf z}}_{1}^{T},\bar{{\bf z}}_{2}^{T},\dots,\bar{{\bf z}}_{N_{\rm RF}}^{T}]^{T}$. Hence we have $\bar{{\bf y}}\sim{\mathcal{C}N}(\mbox{$\mu$},\frac{\sigma^{2}}{L}{\bf I})$ where $\mbox{$\mu$}={\bf E}\mbox{$\omega$}$, and the Fisher Information Matrix (FIM) can be expressed as ${\bf F}_{\mbox{$\eta$}}=\frac{2L}{\sigma^{2}}{\rm Re}\left\{\frac{\partial\mbox{$\mu$}^{H}(\mbox{$\eta$})}{\partial\mbox{$\eta$}}\frac{\partial\mbox{$\mu$}(\mbox{$\eta$})}{\partial\mbox{$\eta$}}\right\}$. Then we have ${\bf C}_{\mbox{$\eta$}}={\bf F}_{\mbox{$\eta$}}^{-1}$, of which the first $M_{t}N_{\rm RF}-1$ diagonal elements represent the corresponding CRBs of estimated variables in $\bf p$}.
@@ IV-B Beam Pattern Optimization
where ${\bf Q}_{n}=\beta\sum_{k=1}^{K}\mathop{\rm diag}({\bf f}^{*}_{k,n}){\bf H}^{H}{\bf w}_{k}{\bf w}_{k}^{H}{\bf H}\mathop{\rm diag}({\bf f}_{k,n})$. Owing to the phase ambiguity removal in Section III-C, we can set $\theta_{t}=0$ and $\phi_{t}=\frac{\pi}{2}$ and thus ${\bf a}_{\rm UPA}(\theta_{t},\phi_{t})={\bf 1}_{M_{t}\times 1}$, where ${\bf 1}_{M_{t}\times 1}$ is an $M_{t}$-element column vector with all entries being 11. Then we have ${\bf w}_{k}^{H}{\bf H}\mathop{\rm diag}({\bf f}_{k,n})=\gamma g_{k}{\bf f}_{k,n}^{T}$ where $g_{k}={\bf w}_{k}^{H}{\bf a}_{\rm UPA}(\theta_{r},\phi_{r})$. Thus ${\bf Q}_{n}$ can be simplified as
@@ V Simulation Results
Fig. 2 provides both the convergence performance and the RMSE performance of the propose algorithm given that $\varepsilon=20^{\circ}$ and $N_{x}=16,32$, i.e., $M_{r}=512,1024$. The left plot shows that the proposed algorithm converges in less than 1010 block descent iterations with SNR =10=-10dB, 0dB, 1010dB. In the right plot, the RMSE performance is simulated using the random generated beam patterns (RD) and the optimized beam patterns (OPT) as SNR varies from 20-20dB to 1010dB. The right plot shows that the RMSEs closely approach the corresponding CRBs and the optimized beam patterns can offer more than 44dB gain of SNR as $M_{t}=1024$, verifying the effectiveness of both the proposed OTA calibration algorithm and the beam pattern optimization method.

### v2

@@ III Over-The-Air Phases Calibration Algorithm
In this section, we propose a block coordinate descent algorithm to solve (6): we first initialize $\bf\Omega$ with all entries being $1$, and estimate the channel $\bf H$}, then given the estimated $\bf H$} and update $\bf\Omega$; as the above iteration goes on, the final result can be obtained until (6) converges. Given the estimated $\bf\Omega$ and $\bf H$}, we also point out the phase ambiguity between them, and remove it by pre-assigning values to the corresponding variables.
@@ IV-A Cramér Rao Bounds
Considering the phase ambiguity, we collect the rest phases of elements of $\bf\Omega$ into the vector $\bf p$}, where ${\bf p}=\left[\angle{\hbox{$\bf\Omega$}_{2,1}},\dots,\angle{\hbox{$\bf\Omega$}_{M_{t},1}},\dots,\angle{\hbox{$\bf\Omega$}_{1,2}},\dots,\angle{\hbox{$\bf\Omega$}_{M_{t}-1,2}},\angle{\hbox{$\bf\Omega$}_{M_{t},N_{\rm RF}}},\right$. $\left.\dots,\angle{\hbox{$\bf\Omega$}_{1,N_{\rm RF}}},\dots,\angle{\hbox{$\bf\Omega$}_{M_{t}-1,N_{\rm RF}}},\angle{\hbox{$\bf\Omega$}_{M_{t},N_{\rm RF}}}\right]^{T}$. Hence for the calibration problem, the unknown parameters can be constructed into $\mbox{$\eta$}=\left[{\bf p}^{T},\theta_{r},\phi_{r},{\rm Real}(\gamma),{\rm Imag}(\gamma)\right]^{T}$. To derive the CRB for the phase calibration problem, we have from (17) that $\bar{{\bf y}}={\bf D}{\mbox{$\omega$}}+\bar{{\bf z}}$, where ${\bf D}=\mathop{\rm blkdiag}\{{\bf C}_{1},{\bf C}_{2},\dots,{\bf C}_{N_{\rm RF}}\}$, $\bar{{\bf y}}=[\bar{{\bf y}}_{1}^{T},\bar{{\bf y}}_{2}^{T},\dots,\bar{{\bf y}}_{N_{\rm RF}}^{T}]^{T}$, ${\mbox{$\omega$}}=[\mbox{$\omega$}_{1}^{T},\mbox{$\omega$}_{2}^{T},\dots,\mbox{$\omega$}_{N_{\rm RF}}^{T}]^{T}$, and $\bar{{\bf z}}=[\bar{{\bf z}}_{1}^{T},\bar{{\bf z}}_{2}^{T},\dots,\bar{{\bf z}}_{N_{\rm RF}}^{T}]^{T}$. Hence we have $\bar{{\bf y}}\sim{\mathcal{C}N}(\mbox{$\mu$},\frac{\sigma^{2}}{L}{\bf I})$ where $\mbox{$\mu$}={\bf D}\mbox{$\omega$}$, and the Fisher Information Matrix (FIM) can be expressed as ${\bf F}_{\mbox{$\eta$}}=\frac{2L}{\sigma^{2}}{\rm Re}\left\{\frac{\partial\mbox{$\mu$}^{H}(\mbox{$\eta$})}{\partial\mbox{$\eta$}}\frac{\partial\mbox{$\mu$}(\mbox{$\eta$})}{\partial\mbox{$\eta$}}\right\}$. Then we have ${\bf C}_{\mbox{$\eta$}}={\bf F}_{\mbox{$\eta$}}^{-1}$, of which the first $M_{t}N_{\rm RF}-1$ diagonal elements represent the corresponding CRBs of estimated variables in $\bf p$}.
@@ IV-B Beam Pattern Optimization
where ${\bf Q}_{n}=\beta\sum_{k=1}^{K}\mathop{\rm diag}({\bf f}^{*}_{k,n}){\bf H}^{H}{\bf w}_{k}{\bf w}_{k}^{H}{\bf H}\mathop{\rm diag}({\bf f}_{k,n})$. Owing to the phase ambiguity removal in Section III-C, we can set $\theta_{t}=0$ and $\phi_{t}=\frac{\pi}{2}$ and thus ${\bf a}_{\rm UPA}(\theta_{t},\phi_{t})={\bf 1}_{M_{t}\times 1}$, where ${\bf 1}_{M_{t}\times 1}$ is an $M_{t}$-element column vector with all entries being $1$. Then we have ${\bf w}_{k}^{H}{\bf H}\mathop{\rm diag}({\bf f}_{k,n})=\gamma g_{k}{\bf f}_{k,n}^{T}$ where $g_{k}={\bf w}_{k}^{H}{\bf a}_{\rm UPA}(\theta_{r},\phi_{r})$. Thus ${\bf Q}_{n}$ can be simplified as
@@ V Simulation Results
Fig. 2 provides both the convergence performance and the RMSE performance of the propose algorithm given that $\varepsilon=20^{\circ}$ and $N_{x}=16,32$, i.e., $M_{r}=512,1024$. The left plot shows that the proposed algorithm converges in less than $10$ block descent iterations with SNR =10=-10dB, 0dB, 1010dB. In the right plot, the RMSE performance is simulated using the random generated beam patterns (RD) and the optimized beam patterns (OPT) as SNR varies from 20-20dB to 1010dB. The right plot shows that the RMSEs closely approach the corresponding CRBs and the optimized beam patterns can offer more than 44dB gain of SNR as $M_{t}=1024$, verifying the effectiveness of both the proposed OTA calibration algorithm and the beam pattern optimization method.